### v4

# Few-Shot Retinomorphic Vision in a Nonlinear Photonic Network Laser

Abstract
With the growing prevalence of AI, demand increases for hardware that mimics the brain’s ability to extract structure from limited data. In the retina, ganglion cells detect features from sparse inputs via lateral inhibition, where neurons antagonistically suppress activity of neighbouring cells. Biological neurons exhibit diverse heterogenous nonlinear responses, linked to robust learning and strong performance in low-data regimes.
Here, we introduce a bio-inspired ‘retinomorphic’ photonic system where spatially-competing lasing modes in a network laser act as heterogeneous, inhibitively-coupled neurons - enabling few-shot classification and segmentation. This compact (150 $\upmu$m) silicon-compatible scheme addresses key challenges in photonic computing: physical nonlinearity and spatial footprint. We report 98.05% and 87.85% accuracy on MNIST and Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis - outperforming software CNNs including EfficientNet in few-shot and class-imbalanced regimes. We demonstrate combined segmentation and classification on the HAM10k skin lesion dataset, achieving DICE and Jaccard scores of 84.49% and 74.80%. These results establish a new class of nonlinear photonic hardware for versatile, data-efficient neuromorphic computing.
Introduction
Artificial intelligence (AI) is increasingly ubiquitous across society and industry. Demand grows for dedicated AI hardware capable of strong performance in use cases where training data and energy are scarce, such as edge computing applications where limited data is collected locally and the power and cooling overheads of GPUs are unattractive. These tasks, including image classification, medical diagnosis, and segmentation, require a range of capabilities:
  • Strong nonlinearity – crucial for harder tasks[1, 2, 3].
  • Extraction of a range of informative features from input data – neural networks capable of this such as convolutional neural networks (CNNs) [4, 5] and transformers[6, 7] typically outperform those without, e.g. multilayer perceptrons (MLPs) [4, 5, 8, 9].
  • The ability to rapidly learn from few training examples – termed the few-shot/low-data regime [10, 11].
Physics-based neuromorphic computing systems[1, 12, 3, 13, 14, 15, 16] have been explored as candidates for edge-AI hardware[14], but to date demonstrations integrating the above qualities and providing software-equivalent performance on harder problems such as biomedical tasks have remained elusive.
Biological systems offer a compelling blueprint. In the retina, ganglion cells perform spatially-sensitive nonlinear processing via lateral inhibition[17, 18, 19, 20], where neurons compete to fire while inhibitively suppressing activity of neighbouring cells. This mechanism enables efficient encoding of spatial structure, even under sparse or noisy input, and the ability to detect many features in parallel from multiple ganglion cells[17, 21]. The combination of excitatory dynamics - firing in response to input, and inhibitory ones - suppressing activity of nearby coupled neurons, dynamics is crucial to the strong performance of biological learning systems, but to date physics-based neuromorphic schemes typically focus solely on engineering excitatory dynamics. Addressing this is an inviting route towards taking neuromorphic performance closer to the biological systems which inspire it. In addition, the diverse heterogenous nonlinear responses across ensembles of biological neurons have been linked to robust learning[22], and creating effective high-dimensional latent spaces which can give rise to strong few-shot learning[11].
Among photonic systems, complex random lasers hosting many interacting modes[23, 24, 25, 26] offer promising neuromorphic platforms. Their lasing thresholds, multiple diverse modes, and competition for optical gain between modes[27, 26, 28, 29] provide intrinsic heterogenous nonlinearity, missing or weak in many photonic schemes[3, 30, 13, 2], while their sensitivity to structured optical input provides spatial processing[27, 29, 26, 31]. The mode competition inhibitively couples modes, as they antagonistically try to harvest a finite amount of pump energy. Despite these properties, such systems have remained unexplored for computation.
Here, we present a retinomorphic machine vision system implemented in a 150 $\upmu$m diameter on-chip semiconductor random network laser[32], fabricated through the mature process of wafer-bonding InP on oxide-coated Si, followed by electron beam lithography and reactive ion etching[33]. The large number of spatially-overlapping lasing modes provide effective heterogenous photonic neurons, performing parallel image feature detection through their spatially-sensitive response, with nonlinearity arising from competition for optical gain between neighbouring coupled modes - a photonic analogue of retinal inhibition. Beyond feature extraction, we demonstrate higher-level vision tasks of classification and segmentation, with strong performance in few-shot/low-data regimes. Crucially, the system outperforms a range of software neural networks including EfficientNet, a large-scale modern CNN, on hard biomedical tasks with scarce and class-imbalanced training data.
These results position the system not just as a novel photonic device addressing the size and nonlinearity issues facing optical computing, but as an exemplar of a new class of compact, few-shot capable neuromorphic hardware. By embedding bio-inspired behaviour combining excitation, inhibition, and nonlinear neuronal heterogeneity directly into the physical dynamics, we enable neuromorphic processing that meets the demands of edge AI.
Figure 1: Working principle of neuromorphic network laser vision scheme. a) Digital micromirror device spatially structures 633 nm, 200 fs pulse pump light into an arbitrary input image, which illuminates a 150 $\upmu$m InP network b). The pump light triggers optical gain in illuminated network regions, exciting photons which scatter through the network to produce many lasing modes at different wavelengths determined by the scattering paths. Spatial profiles of three lasing modes are indicated via different colour network nodes. c) Lasing spectra emitted by the network are recorded via spatially-resolved spectrometer, giving ‘hyperspectral’ data with spectral and spatial dimensions, at $\sim$1 $\upmu$m spatial resolution. Spectral regions which detect image features are highlighted by coloured boxes, corresponding to feature maps in i) and j). d) Schematic of feature detection via kernel filters in software convolutional neural networks. Linear kernel filters are raster-scanned over the input image in small ‘windows’, performing sequential multiply-and-accumulate operations to produce feature maps. Each image feature (e.g. left or right edge) is detected by a separate kernel filter sequentially. As the process is linear, subsequent nonlinear activations must be performed on feature maps to enable vision tasks such as image classification. e) Schematic of network laser feature detection. Small windows of input image are illuminated on the network in a raster-scan. Different lasing modes detect image features in parallel, with integrated physical nonlinearity. Experimentally, ten features are concurrently detected. f) Input ‘Cameraman’ image, inset shows example 4$\times$4 image windows (blue boxes) and raster-scan direction. g) Image windows from f) with corresponding network spectral output (2nd col.), LL plots showing the heterogenous nonlinear response of network modes in response to the different input image pump illuminations (3rd col.), and histograms showing mode intensities for six feature-detecting lasing modes (4th col.). Mode amplitudes vary in response to the presence of different image features, enabling feature detection. h) Composite feature map detected by the InP network. Different feature colours correspond to the lasing amplitude of the six lasing modes. i) Feature maps from the six lasing modes which form the composite map (h), highlighting the range of parallel features detected.
Results and Discussion
Working Principle
[…]
Figure 1d) shows a schematic of feature detection in conventional software CNNs. A kernel filter is raster-scanned across small image ‘windows’ of an input image. The kernel values are matrix multiplied against image pixel values, and the results summed to a single number which becomes the pixel value of an output feature map, centered at the image window location. This process is repeated across the input image to build feature maps, with different kernel filters sensitive to different image features e.g. left or right edges. To detect multiple image features, the process is repeated sequentially using different kernels - consuming time and energy. This process is entirely linear, with subsequent nonlinear activations performed at time and energy cost in order to use the feature maps for higher-level AI tasks such as image classification or segmentation. This linear approach is what has so-far been employed in physical convolutional feature detection schemes[34, 35, 36, 37].
Figure 1e) shows a schematic of the network laser feature detection. Image windows are illuminated on the network in a raster-scan fashion similar to CNNs, but here the network simultaneously detects many features in parallel via multiple lasing modes, with integrated physical nonlinearity. One scan produces many feature maps in parallel, ten experimentally (forty in simulation). Our network occupies a compact 150 $\upmu$m footprint, relative to the mm/cm-scale of alternative physical feature-detecting schemes.
Figure 1f-i) show experimental parallel feature detection results for a ‘Cameraman’ test image (f). Example image windows ($\times 4$ pixels, inset of f) are shown in (g) alongside their respective lasing hyperspectra (2nd col.); LL plots (3rd col., network lasing mode intensity vs pump intensity) which show strongly heterogenous nonlinear responses for each mode, and a variation in these nonlinearities depending on the input image; and a histogram of mode intensities (4th col.) for six feature-detecting modes, taken at the pump power indicated by dashed vertical lines in the LL plots. Mode lasing intensity is increased or suppressed by the presence of different image features, due to a combination of correlation between the input image shape and spatial mode profile, and nonlinear mode competition between spatially-overlapping modes. This enables feature maps to be obtained by plotting the lasing amplitude of different modes above the uniform illumination level at different positions in the raster-scan. The mode intensities plotted in the histogram in (g) are used to generate the pixel amplitudes of our feature maps. Figure 1h) shows a composite feature map produced by six lasing modes, with different colours corresponding to the lasing amplitude of six modes indicated in fig. 1c,g-i). Figure 1j) shows separate feature maps for the modes comprising the composite map. Feature maps for all ten feature-detecting modes are shown in supplementary information.
Figure 2: Parallel image feature-detection via network lasing modes. a-f) show simulation results, g-k) are experimental.
a) 50 $\upmu$m network topology used for netSALT simulations a-f). b) Three pump illumination patterns: uniform, horizontal edge (‘right’) and vertical edge (‘top’). c) Simulated network lasing spectra for the three illumination patterns. Three modes A-C are indicated, corresponding to d). d) Simulated lasing intensities of three modes A-C, for the three pump illumination patterns. Mode competition nonlinearly suppresses specific modes below lasing threshold, and amplifies other modes dependent on input pump image. Spatial mode profiles are shown inset in left panel. e) Input image, ‘CL’ for ‘Complex Lasing’. The input image is fractured into 8$\times$8 pixel windows and illuminated onto the network sequentially in a raster-scan fashion, similar to kernel scanning in a convolutional filter process. f) Simulated feature maps for five modes. Pixel brightness in the feature maps corresponds to mode lasing intensity for that position in the window raster-scan. Different modes lase in response to different image features, spatial mode profiles shown inset. A larger set of 28 simulated feature maps and corresponding mode profiles is shown in supplementary information. g) Experimentally measured ‘CL’ feature maps produced from six modes, with different lasing modes responding to specific image features. h) Experimental composite feature map combining modes from g). i) Equivalent 4$\times$4 kernels for the six lasing modes in (g,h), extracted by linear fits to experimental data. Coloured frames correspond to mode feature map colour scheme. j,k) Colour photograph of a mandrill. Image is separated into RGB channels and feature detection is performed at multiple kernel sizes demonstrating the resolution flexibility of our scheme. Composite RGB feature maps are shown in (k) for 4$\times$4, 7$\times$7 and 11$\times$11 kernels illustrating the range of high-to-low spatial frequency features detected as kernel size is increased.
Parallel Feature-Detection via Network Lasing Modes
We can use simulations of the photonic physics to examine the origin of the feature-detection. Here, we employ netSALT[26], a numerical model which solves the nonlinear interaction of optical waves on graph networks within the steady-state ab-initio laser theory (SALT)[38] approximation. The netSALT model includes amplification and loss on graph edges and nonlinear competition between spatially-overlapping lasing modes. Figure 2a) shows a random Voronoi graph network used for the netSALT simulations. Network topology and average edge length (5 $\upmu$m) match the experimentally measured network shown in fig. 1, with a smaller 50 $\upmu$m diameter used in simulation to reduce computational load.
Three illumination patterns are explored (fig. 2b): uniform illumination (‘full’), horizontal edge (‘right’) and vertical edge (‘top’). The resultant simulated lasing spectra are shown in figure 2c), with strong variation in lasing spectra under the different illumination patterns.
NetSALT predicts 239 possible lasing modes for the network considered here, with 172 reaching lasing threshold for at least one of the illumination patterns. The three spectra in fig. 2c) show the sum of all active modes for the different illumination patterns. We focus on three specific modes, labelled A,B,C, identified in figure 2c) by dashed vertical lines. Figure 2d) shows the lasing amplitude of the three modes under ‘full’, ‘right’ and ‘top’ illumination patterns with the spatial mode profiles inset in the left-hand panel. The modes are spatially-distributed across the network and overlap considerably. As multiple modes are active in the same network waveguides, nonlinear mode-competition occurs between the modes as they antagonistically compete for finite pump gain. Under full illumination, mode B shows the highest amplitude, while mode C does not lase as it falls below threshold. For a right-edge illumination, mode B now falls below the lasing threshold and is deactivated, allowing mode C to lase with the highest amplitude as it no longer faces strong mode-competition from the mode B. For a top-edge illumination, mode A exhibits far higher amplitude and mode C now falls below the lasing threshold.
It is important to note that modes A and C exhibit higher lasing amplitude under ‘top’ and ‘right’ edge illuminations, where the network receives less input pump energy than under uniform illumination. This is not possible in a linear system with non-interacting modes, and occurs here as removing pump light from some network regions causes modes residing even partially in those regions to lose the mode-competition interactions, e.g. mode B wins under full illumination and losing under edge illumination. This observed behaviour shows that nonlinear competition for gain between spatially-overlapping modes is providing feature-sensitive lasing dynamics, analogous to the lateral inhibition processes in retinal ganglion cells.
We now consider the netSALT-simulated feature-detection performance of the system. An input image of the letters CL (‘Complex Lasing’) is shown in fig. 2e). The image was fractured into 8x8 pixel windows and illuminated on the network in a raster-scan fashion. The lasing response was simulated for each window illumination, and the mode lasing amplitude plotted at each position in the raster scan to generate separate feature-maps for each mode. Figure 2f) shows feature maps provided by five different modes, with the corresponding spatial mode profiles inset. These plots demonstrate the ability of the network to perform spectrally-multiplexed feature-detection of a range of image features, with the broad set of lasing modes effectively providing parallel convolutional filters. The spatial location of the mode profiles within the network are seen to show correspondence with the detected image features, e.g. modes detecting lower image edges are located towards the bottom of the network (2nd panel), modes detecting left edges are located towards the left of the network (4th panel), and so on. Of the 172 active modes in the simulated network, 40 were identified as exhibiting good feature-detection. A set of 28 selected feature maps and corresponding spatial mode profiles is shown in the supplementary information, alongside the full set of 172 feature maps from all modes in supplementary information. A regular hexagonal network was simulated to compare against the random topology (supplementary information), with far fewer detected features in the hexagonal network. This demonstrates the ability of network topology to control the device functionality, with further explorations into the correspondence between topology and performance providing fruitful ground for future work.
Figure 3: Network Laser Image classification. a) Single-layer classification. Input images are illuminated directly onto the network, hyperspectra are measured, and a single logistic regression step is applied to the spectra to provide classification. Accuracy of 96.03% is observed for MNIST-digits.
b,c) Multi-layer classification. An initial feature detection layer (b) is implemented as described in figure 2, generating 10 feature maps in parallel via the amplitude of ten lasing modes (indicated by coloured boxes on the hyperspectra). A second classification layer (c) is then implemented by illuminating the feature maps sequentially and measuring hyperspectra. The hyperspectra are combined, and a single logistic regression step applied to return classification predictions. Accuracy of 98.05% is observed for MNIST-digits.
d,e,f) Three classification tasks are implemented, with example images for each task showing detected features: (d) MNIST-digits, (e) Fashion-MNIST, and (f) BreaKHis cancer diagnosis (400x magnification).
g,h,i) Classification accuracy vs. training set length, comparing single- and multi-layer photonic network accuracy against software benchmarks: logistic regression, multilayer perceptron (MLP), LeNet-5 convolutional neural network (CNN), a larger CNN (96 filters), and the far larger EfficientNetV2-B0 CNN (3040 filters, in a transfer-learning scheme with pre-trained ImageNet weights). Maximum accuracies for the multi-layer photonic network of Digits: 98.05%, Fashion: 87.85%, and BreaKHis: 90.12% are observed. Zoomed-in regions of the few-shot regime are shown in (g,h), where the photonic network exceeds software benchmark performance.
We now examine experimental feature detection. Figure 2g) shows six experimentally measured feature maps of the ‘CL’ image with colours corresponding to the modes shown in figure 1 and a composite feature map shown in figure 2h). Parallel feature detection of multiple image features is observed experimentally, with close correspondence to the netSALT simulated behaviour. Supplementary figures examine the feature detection performed by ten lasing modes across a range of test images. To compare our physically-detected feature maps against linear matrix kernel filters as used in software CNNs, we extract effective kernel filters for our modes via least-squares fitting between input images and the experimentally measured mode intensity (details in supplementary information). The extracted effective filters are shown in fig. 2i) with further details in supplementary information. Notably, both positive and negative kernel weights are observed - generating bipolar weights is highly challenging in neuromorphic hardware. This is an attractive benefit of the photonic lateral inhibition processes in our system, which enable mode amplitudes to both increase and decrease in response to specific image features. The $R^{2}$ values below each effective kernel indicate how well the linear fitting process captures our experimental results. When performing the same kernel fitting process on arbitrary linear software kernel filters, the process performs perfectly with $R^{2}$ values of 1.0. The kernels extracted from our experimental data have $R^{2}$ values from 0.628-0.946 (see SI), suggesting that aspects of the physical feature detection process cannot be fully captured by linear kernel filters.
We extend experimental demonstration of feature detection to high-resolution colour images, showing a flexible range of kernel sizes using the popular ‘mandrill’ test image (512$\times$512 pixels, not a baboon as often thought!)[39]. Figure 2j) shows the input image, which is separated into RGB channels and illuminated onto the network at a range of kernel sizes to demonstrate the capability of our system to detect features with arbitrary kernel window size. This flexibility is typically not available in the more common matrix-multiplication style feature detection schemes where kernel size is hard-coded at fabrication stage[34]. Figure 2k) shows experimentally detected feature maps for 4$\times$4, 7$\times$7 and 11$\times$11 kernels, with the R/G/B coloured features detected from the corresponding colour channel (the response of all ten feature-detecting modes are summed to produce a single feature map for each colour channel here). The ability to tune feature detection sensitivity to higher or lower spatial frequency is shown in our feature maps produced by different kernel sizes, the 4$\times$4 kernel picks up high-frequency details in the mandrill fur around image corners, whereas the 11$\times$11 kernel exhibits weaker response to these. Increasingly, powerful modern software CNNs such as EfficientNet[40, 41] employ a range of filter sizes, demonstrating the utility and attraction of our ability to freely select filter size. Separate lasing mode feature maps are shown for a wider range of test images in supplementary information.
Network Laser Image Classification
Here we demonstrate the ability of the network laser to classify images. The network can be operated in a single-layer scheme, directly projecting images onto the network (fig. 3a), or a two-layer scheme with an initial feature-detection layer (fig. 3b), followed by projecting the feature-maps onto the network (fig. 3c). In both cases, a logistic regression step follows the final layer of hyperspectral output to generate classification predictions.
We perform classification on three tasks: MNIST digits, Fashion-MNIST, and the challenging low-data BreaKHis breast cancer diagnosis[42] (colour histopathology images, 400$\times$ magnification, 700$\times$460 resolution). Figures 3d,e,f) show example images for each task, with coloured lines illustrating feature maps detected by the network. For MNIST digits and Fashion, we experimentally process 40,000 training images with a test set of 4000 images and training sets of 125 to 36,000 images. For BreaKHis we process 1693 images, with a test set of 250 and training sets of 10 to 1,443 images. BreaKHis contains a roughly 2:1 ratio of malignant to benign images, with this class-imbalance increasing task difficulty.
Single-layer classification (fig. 3a) works by projecting whole images onto the network, measuring the hyperspectrum, then applying logistic regression on the training set hyperspectra to learn a fixed weight (per image class) for each channel of the hyperspectral output. Once the system is trained, unseen images are classified by performing a weighted sum over the hyperspectral channel amplitudes and learned weights for each class. The class with the largest weighted sum is returned as the network prediction. The single-layer scheme obtains accuracy of 96.03% on MNIST digits, 85.18% on Fashion MNIST, and 86.59% on BreaKHis cancer diagnosis, with measurement throughput up to 100 Hz.
Multi-layer classification begins with feature detection (fig. 3b). Ten spectral regions performing feature detection were selected, producing ten feature maps per input image. Regions returning clear edge-detection were hand selected, with future improvement available by training the feature selection process. Three feature maps are shown schematically in figure 3b), with all feature maps for a range of MNIST digits and Fashion-MNIST shown in supplementary information. These feature maps are then re-illuminated on the InP network as whole images and their corresponding hyperspectra recorded, illustrated in figure 3c). The hyperspectral output from the second-layer feature map illuminations are combined in parallel and a single logistic regression step is applied to produce classification results. The multi-layer scheme obtains accuracy of 98.05% on MNIST digits, 87.85% on Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis. Further details of the regression process are given in the Methods.
Figure 4: Biomedical Image Segmentation and Class-Imbalanced Diagnostic Classification on the HAM10,000/ISIC 2018 dataset. a) Example images of the seven skin lesion classes contained in the dataset. Images are complex with different instances of the same class looking very different, hard to diagnose without expert medical training.
b) Pie-chart illustrating the heavy class-imablance in the dataset.
c) Class-balanced accuracy vs. training set length, comparing photonic network against software benchmarks. Test set is fixed at 1015 images. Class-balanced accuracy penalises incorrect classification on rarer classes more heavily, ensuring performance is not dominated by majority classes. Photonic network scores 53.7% max accuracy, beyond software models.
d) Segmentation scores: DICE, Jaccard, and pixel accuracy of photonic network and software benchmarks. e-g) Image segmentation results, comparing photonic network performance against software benchmarks for three example test set images. Ground-truth masks (yellow line) generated by a human medical expert[43] are used for supervised training, and assessment of network performance on test images. Networks generate predicted segmentation masks (blue line) around where they think the lesion boundaries are. Green regions indicate correct segmentation where the network prediction agrees with the ground truth, red regions show combined false positive & negative errors.
Network laser performance relative to software benchmarks is assessed in figures 3g,h,i). Classification accuracy vs. training set length is assessed for single- and multi-layer network laser schemes against a range of software neural networks. Five software approaches are considered: logistic regression on the raw images (a direct control, effectively bypassing our physical system), a multi-layer perceptron (300$\times$100 neurons), a LeNet5 CNN (22 filters), a larger CNN (96 filters), and the much more powerful ‘EfficientNetV2-B0’[40, 41] CNN (3040 filters, 7.9 million parameters) using a transfer learning approach - initialised with pre-trained ImageNet weights, which are then updated during training. Full details of the software networks are provided in the Methods. We also benchmark against a ‘quasi-isomorphic’ software network: an initial convolutional filter layer of 10$\times$ 4$\times$4 filters followed by 30,000 randomly weighted ReLu neurons and a final logistic regression layer, mirroring the structure of our system. Results are shown in supplementary information.
Few-shot/low-data performance The photonic network laser excels when training data is scarce, a regime where conventional AI can struggle. The network outperforms all considered software models on training sets smaller than 5,000 images (digits), 15,000 images (fashion), and across all training lengths on the harder biomedical BreaKHis task — including the powerful EfficientNet CNN. We score 77.7% accuracy on BreaKHis from just 10 training images (5 examples per class), far higher than the software models. This highlights the effective learning provided by the nonlinear photonic dynamics and mode competition of the network laser. Here we compare performance when using only the raw training set. Software approaches can achieve higher scores on the BreaKHis task, but this requires costly extra steps including artificial augmentation of the training set, and extensive fine-tuning of model weights.
The origin of this strong few-shot performance may lie in the system’s high-dimensional, heterogeneous nonlinear responses. Recent theoretical work has shown that if a learning network is sufficiently ‘high dimensional’ (e.g. it has a large number of output features/channels with unique nonlinear responses), then the network may learn to generalise/accurately classify from very few training examples[11]. This behaviour is exhibited in our network: a broad range of unique nonlinear responses is provided by the many lasing modes, shown experimentally in figure 1g). The diverse nonlinearity in the ‘photonic neurons’ (our lasing modes) provide both high dimensionality and neuronal heterogeneity, which has been shown to promote robust learning and stronger task performance in models of biological neurons[22].
Combined Biomedical Image Segmentation and Class-Imbalanced Diagnostic Classification.
The neuromorphic capabilities of the photonic network are not restricted to feature detection and classification. Here, we demonstrate the ability of the network to spatially map regions of interest in biomedical images - drawing masks over skin lesions in colour photographic images from the HAM10000/ISIC 2018 dataset[44, 43] (fig. 4a, 10015 images of seven different classes of skin lesion). This task, termed segmentation, demonstrates that while the hyperspectral output of our network only contains information on the wavelength and one-dimensional spatial position of the lasing response, it is capable of physically encoding and processing complex two-dimensional spatial information - required for segmentation and other spatial processing. The task here is challenging: lesions are non-uniform in shape, a variety of colours and intensities, and many images contain darker non-lesion regions such as hairs, moles, and areas of uneven lighting. We combine two distinct tasks on the HAM10k dataset: class-imbalanced diagnostic classification, separating lesions into seven classes, and segmentation, drawing spatial masks to indicate lesion location. This two-fold processing has potential future impact in automated/semi-automated combined diagnosis and surgery systems. We need only input the HAM10k images to our photonic network once to achieve both tasks, simply retraining our computationally-cheap linear output regression layer to change between diagnosis and segmentation.
Class-imbalanced diagnosis Here we diagnose lesion images into seven classes. This is made highly challenging by the heavy class-imbalance of the dataset - illustrated in figure 4b). Machine learning typically struggles with class imbalance, tending to incorrectly classify rare classes as examples of the more common classes. Overcoming this requires costly techniques such as data augmentation, e.g. generating many distorted copies of training examples and using them to artificially extend the training set. The HAM10000 task is heavily imbalanced, with the ‘NV’ class representing 66.9% of the dataset, and four of the seven classes each representing 5.1% or less. Figure 4c) shows class-balanced classification accuracy against training set length (class-balanced accuracy penalises incorrect guesses on rare classes more heavily), with the photonic network scoring 53.7% max accuracy, outperforming all software benchmarks at all training lengths. This shows our system demonstrates strong advantages on both few-shot learning and heavily-imbalanced datasets - a hard problem.
Biomedical segmentation For segmentation, we use the same hyperspectral responses from the lesion images which we use for classification. Here, rather than using the diagnostic class as training/testing labels, and logistic regression as an output layer - we use the ground-truth segmentation maps as training/testing labels and ridge regression as an output layer (as the segmentation task has continuous output, rather than discrete labels). Otherwise the procedure follows that used for classification. Figure 4d) compares three segmentation metric scores across the test set: DICE (F1 score for segmentation), Jaccard (intersection over union) and pixel accuracy (correct pixels over incorrect). DICE and Jaccard provide best indication of segmentation quality, pixel accuracy scores tend to be distorted by whether the correct ground-truth mask occupies a small or large fraction of the total image. The photonic network scores highest on all metrics, highlighting the ability of the network to implement spatial processing.
Figures 4 e-g) show segmentation results for the photonic network and software benchmarks for three test images. In the top row (e), the photonic network accurately predicts the lesion location, ignoring the non-lesion mole, unlike the CNN and ridge regression software models. In the middle row (f) the photonic network outperforms software on a complex lesion shape with regions of varying intensity, and in the bottom row (g) the photonic network avoids incorrectly classifying the dark corner regions or hairs.
Conclusions
This work introduces a new class of photonic hardware for compact, highly nonlinear neuromorphic processing which excels at few-shot learning — with potential impact on biomedical sensing, edge computing, and beyond. The system performs parallel feature detection, heterogenous nonlinear activation, and strong image classification and segmentation in a single 150 $\upmu$m on-chip InP network, outperforming powerful modern software CNNs such as EfficientNet particularly on short and imbalanced datasets.
We provide routes to solutions for both the physical nonlinearity and large footprint issues facing many existing photonic computing schemes. As the lasing modes responsible for feature detection spatially overlap, we provide compact feature detection of multiple image features in a micron-scale footprint. In contrast, existing physical convolutional hardware schemes typically detect one feature per physical repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected.
Strong learning in the few-shot/low-data regime with scarce training examples is a hard and open challenge in machine learning. The performance of our photonic network in this regime, including on hard biomedical tasks such as BreaKHis and HAM10k, provides valuable evidence that physics-based neuromorphic systems can handle complex, data-constrained learning tasks. These results bring physical neuromorphic schemes closer to offering viable alternatives to conventional software-based approaches. Our good performance here extends to heavily class-imbalanced datasets, which are similarly challenging for conventional software AI. One of the key proposed applications for neuromorphic computing hardware is ‘edge computing’ where remotely deployed hardware performs computation directly at the point of sensing, adapting on-the-fly to incoming data and changing environmental conditions without access to large offline/cloud-based training data. Here, it is highly advantageous to perform well with short and imbalanced training sets, as locally collected data will typically be far from the ideal large and class-balanced training sets where software AI is most comfortable.
A key result of this work is demonstrating the importance of including both excitatory and inhibitory interactions in neuromorphic hardware. In biological systems, both of these interactions and their interplay are crucial to providing neuronal functionality[19, 21, 17], and existing physical neuromorphic hardware platforms often focus solely on excitatory signals. By engineering inhibitory nonlinear mode-competition between network lasing modes, we demonstrate the functional benefits of bio-inspired inhibitory dynamics.
As we develop finer control over the network dynamics, one can envision networks engineered to detect arbitrary specific features or provide reconfigurable detection via altering network dynamics through electrical control or selective illumination of network sub-regions[26] to provide effective tunable physical weights in the input space[45]. More broadly, our approach shows the benefits of designing physical computing systems with diverse physical nonlinearities, and a balance of competing excitatory and inhibitory interactions. While our system currently exhibits strongest performance on harder, lower training data tasks, we foresee that future iterations will extend this performance to larger datasets via training of the physical dynamics and optimisation of the network graph topology.
Author contributions
RS, JCG, KM and KS conceived the work.
RS, KM, JCG, KS and WRB secured funding.
WKN, JD, DS, TF and AF performed optical experiments.
JCG and KS designed the neuromorphic computing architecture.
JCG, JD and TVR developed the feature detection scheme.
RS, KM, HS, J.D., AF, DS, and W.K.N. designed and developed the network laser concept.
JD, KM and HS fabricated the InP network.
TVR performed all simulations.
JCG, JD, KS and TF performed machine learning analysis.
JD and TS performed analysis of experimental feature-detection functionality.
JD selected the biomedical datasets studied.
MB provided guidance on physical simulations using the netSALT code and guidance on implementing the machine learning architecture.
JP designed a benchmarking task used to develop the computing architecture.
AF, WKN and JD designed and produced the figures.
JCG wrote the manuscript with contributions from all authors.
Acknowledgements
This work was supported by the Royal Academy of Engineering Research Fellowships, awarded to JCG.
JCG was supported by the EPSRC ECR International Collaboration Grant EP/Y003276/1.
AF, JD, HM, RS and KM acknowledge support from the EU ITN EID project CORAL (GA no. 859841).
TVR, DS, and RS acknowledge support from The Engineering and Physical Sciences Research Council (EPSRC), grant number EP/T027258.
WKN acknowledges the research support funded by the President’s PhD Scholarships from Imperial College London.
This work was supported by the EPSRC Impact Acceleration Account fund awarded to JCG and RS.
This work was supported by the Imperial College London DT-Prime scheme awared to JCG, RS and MB.
KDS was supported by The Eric and Wendy Schmidt Fellowship Program and the Engineering and Physical Sciences Research Council (Grant No. EP/W524335/1).
TF was supported by the EPSRC doctoral training programme PhD award.
JP performed the work as part of his Physics MSci project at Imperial College London.
The InP network laser samples were fabricated in the Binnig and Rohrer Nanotechnology Center (BRNC) at IBM Research Europe - Zurich. We thank Markus Scherrer, Daniele Caimi and the Cleanroom Operations Team of the BRNC for their help and support.
We thank David Mack for excellent laboratory management and Stephen Cussell and David Bowler for excellent technical workshop services at Imperial College London.
We thank Christine Zhao and Eunju Moon for their helpful discussions and contributions to the project.
Simulations were performed on the Imperial College London Research Computing Service[46].
Test image of ‘Cameraman’ was used with CC BY-NC 4.0 creative commons licence, first appearance in Schreiber et al[47]. Obtained from dome.mit.edu/handle/1721.3/195767.
Test image of ‘Street Scene’ was used with CC0 1.0 universal creative commons licence, photographer Christian Birkholz, taken 2013. Obtained from pixabay.com/photos/calle-jose-betancort-teguise-201343/.
Test image of ‘CL’ letters was generated by the authors.
Test image of ‘Mandrill’ was taken from the USC-SIPI image database, available at: https://sipi.usc.edu/database/database.php
Competing interests
The authors declare no competing interests.
Data availability statement
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
Code availability statement
The code used in this study is available from the corresponding author on reasonable request.
Methods
Experimental methods
Fabrication
The InP networks are fabricated on a Si chip with a 2 $\mu$m buffer layer of SiO2 according to [33]. To obtain a defect-free InP layer on SOI, we used a Direct Wafer Bonding approach where the 127 nm thick InP layer is grown on a sacrificial III-V semiconductor wafer, with lattice-constant matched growth. This wafer is then annealed onto the SOI wafer and the sacrificial wafer removed by wet etching, resulting in a InP on SOI wafer that is cleaved into chips. The network structures are then patterned into a spin-coated hydrogen silsesquioxane (HSQ) mask with electron beam lithography. The pattern is transferred into the InP layer using Inductively Coupled Plasma (ICP) Reactive Ion Etching (RIE)[33] using a cyclic two-stage process: a CH4 and H2 plasma stage for etching, followed by O2 plasma stage for cleaning. To reduce surface state recombination losses and oxidation, the structures are passivated with phosphoric acid and coated with 3 nm of Al2O3 using Atomic Layer Deposition (ALD) [48].
Network design
The network was designed by using a randomly-distributed point grid pattern on a plane and generating a Voronoi diagram[32]. The network is cropped to a circular area with 150 $\upmu$m diameter, and the density of the point patterns is varied to generate waveguide graph-edges with 350 nm width and an average length of 5 $\upmu$m.
Optical setup
[…]
Throughput of images was run up to 100 Hz, with the limiting factor currently the camera detection. We have run experimental tests on a similar system up to 500 Hz. With the correct camera, throughputs of 1 kHz are readily feasible without modification to the scheme. Beyond this, the next limiting factor is the rate at which the DMD can display new images. Our current DMD has 10 kHz speed, and faster DMDs are available up into the MHz, with ongoing improvements in DMD technology likely to improve this in the near future. The repetition rate of our pulsed laser goes to 100 kHz, with MHz pulsed lasers or on-chip light sources available. The eventual fundamental limit is the 10-100 ps timescale of the photonic dynamics. Beyond this, more fundamental changes to the physical system/excitation scheme will be required.
A second 80 picosecond pulse pump laser was used to confirm that we are not limited to operating with femtosecond pulses, and can operate with a range of pump wavelengths. We employed the Picophotonics CP32 532 nm MOPA pump laser, which successfully pumps our InP network to lase. Experimental data confirming this (hyperspectra, LL plots) are shown in the supplementary information.
Hyperspectral measurements
The raw hyperspectral data recorded by the CCD has 1024 spectral channels and 256 spatial channels. For feature detection, hyperspectral regions comprising four adjacent spectral channels and eight adjacent spatial channels are used to generate each feature map, with the averaged intensity taken within the hyperspectral region.
For the image classification layer, the hyperspectra are divided into averaged regions of sixteen adjacent spectral channels and four adjacent spatial channels, with each region passed to the logistic regression layer and assigned a single fixed linear weight during training.
The region-selection and averaging performed here could be implemented by using a lower resolution spectrometer, with the additional benefits of faster data acquisition and lower cost.
NetSALT simulations
The lasing modes and intensities of the network are modelled using netSALT[26]. The designed network is represented as a graph, with the edges subdivided to overlap with an 8$\times$8 grid to correspond to the kernel. When the network is unpumped, each edge has a complex refractive index n = 3.4 + 0.001i. The edges are further assumed to have a Lorentzian gain profile centered at normalised frequency $k_{a}=7$ with a width $\gamma_{a}=0.35$, and are pumped up to a normalised power $D_{0}=0.5$ (For a detailed physical explanation of the parameters, see Ref. 26). A separate simulation is performed for each 8$\times$8 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane $k$, and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im$(k)=0$], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.
As netSALT provides one mode frequencies and intensities but not the widths, the realistic spectra in Fig. 2c are created by assigning a width of 0.005 to each mode and adding a broad photoluminescence background. The mode feature maps in Fig. 2f are made by selecting specific modes and plotting their intensities for the kernels corresponding to the kernel scan of Fig. 2e.
Regression process for the network laser image classification task
Hyperspectra are recorded for the ten feature-maps for each of the 40,000 input images, with initial hyperspectral resolution of 1024 spectral channels and 256 spatial channels. The top and bottom 25 spatial channels are discarded as they fall outside the InP network, leaving 206 spatial channels.
These channels are then binned using simple averaging with binning factors of 16 and 2 for the spectral and spatial channels to give 64 and 103 spectral and spatial channels respectively per feature map.
The spectral and spatial channels of all feature maps are then combined in parallel to produce a single 2D array, with each row of the array corresponding to a single input image and each column corresponding to a single ‘pixel’ of the binned hyperspectra.
The array is separated into separate ‘train’ and ‘test’ sections, with final accuracy results of 98.05% (MNIST-digits) and 87.85% (Fashion-MNIST) reported from train and test lengths of 36,000 and 4000 images respectively.
Feature selection is performed on the data using the SelectKBest function from Sklearn, using Chi-squared correlation between the hyperspectral data and the ground-truth image class labels in the training set. A parameter search was performed to identify 36,000 as the number of features to keep from the initial pool of 58,000 features.
After feature selection, logistic regression is performed on the training data using sklearn’s logisticRegression function. The lbfgs solver was used with a regularisation parameter of C = 6.8, identified using a parameter search.
Regression process for the network laser image segmentation task
The process is nominally identical to the image classification workflow, save for these details on replacing logistic regression with ridge regression, and changing the ground truth from classification labels to segmentation maps:
Ridge regression is used instead of logistic regression (using sklearn’s built in ridge regression function) as we want continuous outputs here. Logistic regression returns discrete outputs, such as class-based classification predictions. Segmentation is performed by generating 2D maps of confidence/probability that a pixel is part of a lesion, hence ridge regression is used as the output layer. The confidence maps are then thresholded at 50% confidence to produce binary maps, and regions with higher than 50% are predicted to contain lesions. We experimented with varying thresholds and having this as an optimised hyperparameter, but the best performing threshold values were so close to 50% (plus or minus 1% typically) that we leave the threshold fixed at 50% to simplify the process.
The ground truth labels used for training and testing in the segmentation tasks are 2D maps produced by a human medical expert[43]. We train fixed weights on each hyperspectral output channel which generate segmentation maps based on network hyperspectra, trained by maximising DICE accuracy against the ground truth maps. We train on 9000 images, and test on 1015.
Software machine learning comparisons for image classification
We benchmark our system against five software approaches for the classification tasks: logistic regression, a multi-layer perceptron, a small convolutional neural network (LeNet-5), and a larger convolutional neural network model with more filters (96) and trainable parameters, and a far larger EfficientNetV2-B0 CNN (7.9 million parameters, 3040 filters) pre-trained using ImageNet weights and operated in a transfer learning scheme.
The model descriptions and number of trainable parameters for each model are given below, models were coded in python using the sklearn, keras and pytorch packages.
Multi-layer Perceptron
The multi-layer perceptron comprises an input layer of 784 neurons (one per 28x28 image pixel), an initial hidden layer of 300 neurons with ReLU nonlinear activation, a second hidden layer of 100 neurons with ReLU nonlinear activation, and then a final softmax layer.
  • Input Layer:
    • Neurons: 784 (one per 28x28 image pixel)
  • Hidden Layer 1:
    • Neurons: 300
    • Activation: ReLU (Rectified Linear Unit)
  • Hidden Layer 2:
    • Neurons: 100
    • Activation: ReLU
  • Output Layer:
    • Neurons: 10
    • Activation: Softmax
Number of Trainable Parameters:
[…]
LeNet-5 CNN Structure:
The LeNet-5 convolutional neural network architecture comprises seven layers: two convolutional layers each followed by a max pooling layer, two fully connected layers then a final softmax output layer. In total, 22 convolutional filters are used.
  • Convolutional Layer 1:
    • Filters: 6
    • Kernel Size: $5\times 5$
    • Activation: ReLU
    • Input Shape: $28\times 28\times 1$
  • Max Pooling Layer 1:
    • Pool Size: $2\times 2$
    • Strides: $2\times 2$
  • Convolutional Layer 2:
    • Filters: 16
    • Kernel Size: $5\times 5$
    • Activation: ReLU
  • Max Pooling Layer 2:
    • Pool Size: $2\times 2$
    • Strides: $2\times 2$
  • Flatten Layer
  • Fully Connected Layer 1:
    • Neurons: 120
    • Activation: ReLU
  • Fully Connected Layer 2:
    • Neurons: 84
    • Activation: ReLU
  • Output Layer:
    • Neurons: 10
    • Activation: Softmax
Number of Trainable Parameters:
[…]
Total Trainable Parameters: $780+4,832+48,120+10,164+850=64,746\text{ parameters}$
Larger CNN model (96 filters)
The larger CNN model has a more complex structure than the LeNet-5 CNN, with more kernel filters and trainable parameters. In total, 96 convolutional filters are used.
  • Convolutional layer 1:
    • Input channels: 1
    • Number of filters: 32
    • Kernel size: $3\times 3$
  • Max pooling layer 1:
    • Kernel size: $2\times 2$
  • Convolutional layer 2:
    • Input feature maps: 32
    • Number of filters: 64
    • Kernel size: $3\times 3$
  • Max pooling layer 2:
    • Kernel size: $2\times 2$
  • Fully connected layer 1:
    • Input size: $64\times 5\times 5$ = 1600
    • Output size: 128
  • Fully connected layer 2:
    • Input size: 128
    • Output size: 10
Number of Trainable Parameters
The number of trainable parameters in the model is calculated as follows:
Conv1 parameters $=(1\times 32\times 3\times 3)+32=320$
Conv2 parameters $=(32\times 64\times 3\times 3)+64=18,496$
FC1 parameters $=(64\times 5\times 5\times 128)+128=204,928$
FC2 parameters $=(128\times 10)+10=1,290$
Total Trainable Parameters $=320+18,496+204,928+1,290=225,034$
EfficientNetV2-B0 Transfer Learning Model
This model uses transfer learning using the EfficientNetV2-B0 architecture with an added classification head MLP. EfficientNetV2 is a relatively modern (2019[40] original EfficientNet model, 2021[41] for the EfficientNetV2 employed here) large convolutional neural network, frequently deployed in industrial and commercial learning applications and applied to real world problems. EfficientNetV2-B0 has 3040 convolutional filters, far more powerful than the other CNNs considered here. As is common in learning applications without sufficiently large training sets to fully optimise network parameters from scratch, pretrained weights learned from training on the ImageNet dataset are used to initialise the network and set up a broad range of feature extracting kernels before the network is optimised on the training set of an arbitrary desired task.
The EfficientNetV2-B0 base model is followed by a 512 $\times$ 128 neuron multilayer perceptron classification head to aid repurposing of the initialised ImageNet weights to arbitrary tasks, comprising: global average pooling, two fully connected (dense) layers with ReLU activations, and a final softmax classification layer.
  • EfficientNetV2-B0 (pretrained on ImageNet) with 512 $\times$ 128 MLP head
    • Stem Convolution Layer:
        Filters: 32
        Kernel Size: $3\times 3$
        Stride: 2
        Activation: Swish
    • Fused-MBConv Blocks:
        Stage 1: 1 block, 16 filters, expansion: 1, kernel size: $3\times 3$
        Stage 2: 2 blocks, 24 filters, expansion: 4, kernel size: $3\times 3$
        Stage 3: 2 blocks, 32 filters, expansion: 4, kernel size: $3\times 3$
    • MBConv Blocks:
        Stage 4: 3 blocks, 48 filters, expansion: 4, kernel size: $3\times 3$
        Stage 5: 5 blocks, 64 filters, expansion: 6, kernel size: $3\times 3$
        Stage 6: 8 blocks, 128 filters, expansion: 6, kernel size: $3\times 3$
        Stage 7: 1 block, 160 filters, expansion: 6, kernel size: $3\times 3$
    • Head Convolution Layer:
        Filters: 1280
        Kernel Size: $1\times 1$
        Activation: Swish
    • Global Average Pooling:
        Reduces spatial dimensions to a single 1280-dimensional feature vector
    • Dense Layer 1:
        Neurons: 512
        Activation: ReLU
    • Dense Layer 2:
        Neurons: 128
        Activation: ReLU
Number of Trainable Parameters:
EfficientNetV2-B0 Base: $\approx 7.2\text{ million parameters}$
Dense Layer 1: $(1280+1)\times 512=655,872$
Dense Layer 2: $(512+1)\times 128=65,664$
Output Layer: $(128+1)\times 10=1,290$
Total: $\approx 7,922,826\text{ parameters}$